OpenStudy (anonymous):

what is the solution of square root of x+2

OpenStudy (anonymous):

What do u mean?

OpenStudy (anonymous):

Is it \[\sqrt{x}=2\]

OpenStudy (anonymous):

\[\sqrt{x +2} = 0 \implies x = -2 \]